# Semantic Segmentation , Self-Driving Car , Computer vision , BDD100K , Light-weight model , Embedded devices

### v4

@@
Semantic segmentation is crucial for autonomous driving, particularly for the tasks of Drivable Area and Lane Segmentation, ensuring safety and navigation. To address the high computational costs of current state-of-the-art (SOTA) models, this paper introduces TwinLiteNetPlus (TwinLiteNet+), a model capable of balancing efficiency and accuracy. TwinLiteNet+ incorporates standard and depth-wise separable dilated convolutions, reducing complexity while maintaining high accuracy. It is available in four configurations, from the robust 1.94 million-parameter TwinLiteNet to the ultra-lightweight 34K-parameter TwinLiteNet. Notably, TwinLiteNet attains a 92.9% mIoU (mean Intersection over Union) for Drivable Area Segmentation and a 34.2% IoU (Intersection over Union) for Lane Segmentation. These results achieve remarkable performance, surpassing current state-of-the-art models while only requiring 11 times fewer Floating Point Operations (FLOPs) for computation. Rigorously evaluated on various embedded devices, TwinLiteNet+ demonstrates promising latency and power efficiency, underscoring its potential for real-world autonomous vehicle applications. The code is available on https://github.com/chequanghuy/TwinLiteNetPlus.
@@
Semantic Segmentation , Self-Driving Car , Computer vision , BDD100K , Light-weight model , Embedded devices
††journal:
[…]
top=1.1in, bottom=1.5in, left=0.7in, right=0.7in
1 Introduction
The emergence of deep learning methodologies has driven significant growth in the field of autonomous vehicles, making it a key area of research within artificial intelligence and computer vision. In the context of autonomous navigation, the efficacy of the vehicle’s decision-making mechanisms is critically dependent upon the precision with which the system can identify and comprehend its surroundings. Self-driving cars often use sensors such as Radar, LIDAR or cameras to perceive their environment during movement. While these technologies function across diverse weather conditions and provide depth information from the environment, Radar and LIDAR are more expensive in comparison to cameras and, notably, cannot discern colors. This limitation leads to an increase in the overall cost of the sensor systems for autonomous vehicles without offering the detailed environmental imagery provided by cameras. Consequently, within the domain of practical applications for autonomous vehicles,
Figure 1:
Comparison of evaluation metrics mIoU (%) (for Drivable Area Segmentation) - IoU (%) (for Lane Segmentation) - GFLOPs of various models on the BDD100K dataset. The x-axis represents the number of GFLOPs of the model, the y-axis represents the mIoU (%) for the Drivable Area Segmentation task, and the radius of the circle indicates the IoU (%) for the Lane Segmentation task (the larger the radius, the higher the accuracy).
cameras have remained a principal focus and have undergone rigorous development, particularly in scenarios where deep learning is extensively employed for image analysis and, more broadly, for advancements in computer vision.
In recent years, techniques related to transformers and convolution-based methods have undergone extensive research in the field of computer vision, particularly for specific tasks such as classification, object detection, and segmentation. While transformers have made significant progress in natural language processing and image analysis tasks, they often result in high latency and require a substantial number of parameters and extensive data to achieve optimal results. For autonomous vehicle models, ensuring low latency is essential, even if it comes at a slight cost to accuracy. Therefore, in the context of autonomous driving, research continues to place a strong emphasis on convolution-based models.
In Advanced Driver Assistance Systems (ADAS), semantic segmentation plays a pivotal role in classifying drivable areas and lane markings. This capability significantly enhances an autonomous vehicle’s navigation and obstacle avoidance, thereby ensuring safer operation. The accurate detection of lane markings and the differentiation between directly drivable lanes and alternate lanes are crucial for facilitating precise steering and lane-change decisions. Although research has produced many models for segmentation tasks within the autonomous driving context, such as drivable area segmentation [1, 2] or lane segmentation [3, 4, 5], achieving promising results, a naive implementation approach would lead to a linear increase in the number of models matching with the task complexity. To optimize both accuracy and inference time, recent advancements have favored the adoption of multi-task models over single-task models.
Multi-task models, capable of simultaneously executing drivable area segmentation and lane segmentation [6, 7, 8], alongside incorporating object detection functionalities [9, 10, 11], are gaining significant interest due to their efficiency in handling multiple tasks concurrently. However, recent developments in these models mainly concentrate on enhancing model accuracy, often overlooking the practical applicability of such models in autonomous vehicle applications. Multi-task models [7][9] are constructed with considerable complexity to facilitate the simultaneous execution of numerous tasks. These models are typically evaluated on high-end hardware, such as RTX3090 and Tesla V100, achieving processing speeds of several tens of frames per second. Nevertheless, these evaluations usually neglect considerations when deploying on embedded devices, which involve in inference speed and power consumption. Focusing only on accuracy makes it difficult to use these complex models in self-driving cars, especially those that have limited computing power. Large-scale models or those not initially designed for deployment on embedded devices often utilize model compression techniques such as Pruning, 8-bit Quantization, and Knowledge Distillation for implementation on embedded systems. These methods, however, can cause accuracy degradation compared to the original model, and not every model successfully maintains high accuracy after applying these techniques. To tackle this challenge, this paper introduces a multi-task model suited for segmentation tasks within the autonomous driving context, designed with an emphasis on efficient execution on devices characterized by limited computational capacity, specifically embedded devices.
@@ 1 Introduction
In this work, we introduce TwinLiteNet+, a model tailored for real-time operation with optimal power consumption and hardware resources, proficient in simultaneously segmenting lanes and drivable areas. This model demonstrates competitive precision compared to models with similar tasks. There are three principal contributions of our research: (1) We present a lightweight Convolutional Neural Network (CNN) model, optimized for low computational cost, consisting of a singular encoder block and dual decoder blocks for drivable area and lane segmentation respectively. (2) The model is available in four distinct configurations, each optimizing the trade-off between accuracy and computational efficiency. (3) To validate our model’s applicability in real-world conditions, we deploy and assess its performance in terms of speed and energy efficiency on a range of embedded devices, including the Jetson Xavier and Jetson TX2. The remainder of the paper is represented as follows: We evaluate relevant models in Section 2 to grasp the benefits and drawbacks in the tasks of Drivable Area Segmentation, and Lane Segmentation with Multi-task approaches. The proposed TwinLiteNet+ presents an architecture with methods to boost model performance in Section 3. In Section 4, we conduct experiments on the BDD100K dataset, and the results show that our TwinLiteNet+ models outperform on latency and power efficiency. In the final Section, we provide some conclusions and future development directions.
2 Related works
In this section, we present a brief evaluation of related work, focusing particularly on segmentation models in the context of autonomous vehicles and methods for optimizing deep learning models on embedded devices.
2.1 Semantic Segmentation
Semantic segmentation is a prominent research area extensively explored in the field of computer vision. The main distinction separating it from detection tasks lies in its ability to perform pixel-level classification, involving the labeling of each pixel to delineate objects and their boundaries. Initially, Convolutional Neural Network (CNN) lays the groundwork for effective methods in handling these tasks, with many high-performance models developed for numerous applications [12, 13, 14]. Furthermore, the integration of attention modules [15][16] has opened up powerful approaches for semantic segmentation, significantly enhancing the learning capabilities of segmentation models during training.
2.1.1 Drivable Area Segmentation
Recent works have suggested many efficient approaches for semantic segmentation in self-driving tasks, particularly in drivable area segmentation, surpassing traditional lane segmentation models that can only recognize the road ahead of a vehicle. Both [17, 18] employ deep learning methods to predict drivable area utilizing a ResNet backbone. While the first study enhances semantic information using Feature Pyramid Network (FPN) and Atrous Spatial Pyramid Pooling (ASPP) modules, the second integrates LinkNet [19] and a CycleGAN-based augmentation to improve night-time segmentation. Zhao et al.[1] devise the PSPNet model utilizing the Pyramid Pooling Module (PPM) that applies global average pooling with multiple different bin scales to extract various levels of features for dividing drivable area. To further enhance applicability, ESPNet[13] uses Dilated Convolutions to build an efficient spatial pyramid (ESP) module that can deliver real-time processing and still maintain high accuracy compared to the aforementioned models.
2.1.2 Lane Segmentation
Within the field of autonomous vehicles, lane detection is also among the basic topics for driving perception. Old-fashioned lane-detecting methods have been conducted such as using a color-based method to extract lane boundary [20], or converting to HSI Color [21] to detect lane markings based on thresholds. These conventional methods heavily depend on handcrafted rules or thresholds and often struggle with environmental conditions such as lighting or weather. They require domain insights for parameter tuning and are often unable to adapt to diverse real-road conditions. To solve this, Global Association Network (GANet) [22] is proposed to handle this task from a keypoint-based perspective or [23] introduces a transformer-only method in its framework to enhance lane segmentation through local prior knowledge. In recent years, segmentation-based methods in deep learning have proven to be the most practical approach for many studies. In 2017, Angshuman Parashar et al.[3] proposed Spatial CNN (SCNN) with a new neural network architecture by generalizing traditional deep layer-by-layer to slice-by-slice convolution in feature mapping, allowing information to propagate across pixels along rows and columns within a layer despite time-consuming. On the other hand, Enet-SAD [4] utilizes a self-attention-guided filter method to assist low-level feature maps in learning from high-level feature maps, or recently, road marking methods [24, 25] has also brought much attention from the community. Later on, a novel detector CLRerNet [26] along with a proposed LaneIoU metric are presented to further improve the confidence score. Although previous research highlighted the initial success of models in the single-task domain, the advantage of leveraging a model to address multiple tasks has led to recent studies demonstrating that a multi-task strategy is more appropriate for real-world applications, particularly when deployed on devices with limited computing capabilities, such as embedded systems. This method allows models to strike a balance between efficiency and complexity while concurrently performing two tasks: segmentation of the driver’s zone and lane detection. Consequently, we introduce an architecture designed with shared encoder blocks that distribute information to various decoder blocks.
Figure 2: The TwinLiteNet+ architecture comprises two phases. During the Encode phase, the input image passes through an Encoder block followed by a Partial Class Activation Attention mechanism. In the Decode phase, the output from the Encoder is channeled through two identical yet independent Decoder blocks, transforming the feature maps into two separate segmentation maps.
@@ 2.2 Multi-task Approaches
Considering the shift from single-task models to multi-task frameworks, this approach has become an established strategy for concurrently addressing multiple tasks. It enhances shared representations and leverages the similarities among diverse tasks. The introduction of the BDD100K dataset has propelled research into multi-task models for autonomous driving challenges, leading to the exploration of various models, including those for drivable area segmentation and lane segmentation [8, 6, 7]; drivable area segmentation and scene classification [27, 28]; as well as drivable area segmentation, lane segmentation, and object detection [10, 29, 11]. [10, 29, 11] introduce a model that integrates a YOLO-based shared backbone with an encoder-decoder structure, effectively combining three separate sensory tasks: vehicle detection, drivable area segmentation, and lane detection. A comparable encoder-decoder architecture is also present in Hybridnets [9], which features a more lightweight backbone through the use of depth-wise separable convolutions. Recently, CenterPNets [30] has gained attention for its ability to achieve high accuracy and precision with an end-to-end shared multi-task network. However, the previously proposed multi-task models were primarily focused on improving accuracy and were not extensively tested on devices with limited computational capabilities. As a result, directly implementing these models in autonomous vehicle systems continues to pose significant challenges.
Contrary to the goals of previous approaches, TwinLiteNet [8] is specifically designed to facilitate real-time implementation on devices with low configurations, particularly embedded systems. With merely 0.44M parameters, this model presents competitive accuracy in tasks such as lane segmentation and identifying drivable areas. Significantly, it incorporates an ESP encoder, which is based on dilated convolutions, enabling efficient feature extraction with a reduced number of parameters. Despite TwinLiteNet’s efficient streamlined architecture, the simplistic design of its decoder block limits the model’s ability to effectively leverage information during the decoding process. Inspired by TwinLiteNet, our TwinLiteNet+ model improves upon the encoder by integrating additional convolutional layers after the Transpose Convolution layer, thereby enhancing feature extraction while still maintaining low computational costs. Moreover, the introduction of skip connection techniques during the decoding phase enables the combination of input and decoding features. Additionally, in Encoder block of TwinLiteNet+, Depthwise and dilated convolutions are combined to robustly optimize the model’s latency.
@@ 2.3 Optimizing Deep Learning Deployment on Embedded Devices
In recent years, deploying deep learning models on embedded devices has attracted considerable attention within the artificial intelligence community. The computation on these devices is distinguished by its ability to execute processes directly on units that are both cost-effective and task-specific. Despite this advantage, such capabilities are constrained by significant limitations in terms of computational power and storage capacity, presenting substantial challenges for the deployment of deep learning models on these platforms. Different model compression techniques, such as 8-bit quantization, pruning, and knowledge distillation, have been devised to enable lower-cost computation without sacrificing accuracy. Quantization, in particular, is a straightforward yet efficacious method that involves performing calculations with fewer bits than the standard 32-bit representation. Post-training quantization enables models, initially trained with 32 bits, to conduct inference at reduced bit levels without necessitating retraining. In contrast, Quantization-aware training incorporates Fake Quantize layers, usually resulting in greater accuracy compared to post-training quantization. However, the latter is preferred for its simplicity, as it obviates the need for retraining the model. Howard et al. introduce MobileNet[31], a class of efficient models tailored for embedded and mobile vision applications, employing depthwise separable convolutions to mitigate computational demands. Recent studies have demonstrated the superiority of Depth-wise dilated separable convolutions over standard Dilation in classification tasks, achieving an accuracy of 67.9% with 123M FLOPs, as opposed to 69.2% accuracy at the cost of 478M FLOPs (an increase of 3.9) with standard Dilation. This finding underscores that, although standard Dilation achieves marginally higher accuracy, it does so at the expense of a considerable increase in computational cost. Furthermore, the development of frameworks such as TensorFlow Lite, TensorRT, ncnn, and MNN[32] has been instrumental in streamlining the deployment of deep learning models on embedded devices. Despite the availability of numerous techniques to facilitate model deployment on devices with limited computational capacity, achieving a balance between computational cost and latency remains crucial for real-time inference. This balance is particularly important as models often prioritize accuracy at the expense of computational complexity. In this research, we introduce an efficient computational model that is optimized for resource-constrained devices without depending on any model compression techniques. Our proposed model combines the robustness of dilated convolutions with the rapidity of depthwise separable convolutions within the Encoder block. This approach effectively ensures both high accuracy and low latency, optimizing the model’s performance for practical applications.
Algorithm 1 Feedforward process of TwinLiteNet+
1:The input image I and two downsampled image I1 and I2I1 and I2 are downsampling images derived from equations 3 and 4, respectively.
2:Drivable Area Segmentation Map and Lane Segmentation Map
3:F Encoder(I, I1, I2)
4:F PCAA(F)
5:O Decoder(I, I1, I2)
6:O Decoder(I, I1, I2)
7:returnO, O

### v5

# TwinLiteNet+: An Enhanced Multi-Task Segmentation Model for Autonomous Driving

@@
Abstract
Semantic segmentation is a fundamental perception task in autonomous driving, particularly for identifying drivable areas and lane markings to enable safe navigation. However, most state-of-the-art (SOTA) models are computationally intensive and unsuitable for real-time deployment on resource-constrained embedded devices. In this paper, we introduce TwinLiteNet+, an enhanced multi-task segmentation model designed for real-time drivable area and lane segmentation with high efficiency. TwinLiteNet+ employs a hybrid encoder architecture that integrates stride-based dilated convolutions and depthwise separable dilated convolutions, balancing representational capacity and computational cost. To improve task-specific decoding, we propose two lightweight upsampling modules-Upper Convolution Block (UCB) and Upper Simple Block (USB)-alongside a Partial Class Activation Attention (PCAA) mechanism that enhances segmentation precision. The model is available in four configurations, ranging from the ultra-compact TwinLiteNet (34K parameters) to the high-performance TwinLiteNet (1.94M parameters). On the BDD100K [1] dataset, TwinLiteNet achieves 92.9% mIoU for drivable area segmentation and 34.2% IoU for lane segmentation-surpassing existing state-of-the-art models while requiring 11× fewer floating-point operations (FLOPs) for computation. Extensive evaluations on embedded devices demonstrate superior inference speed, quantization robustness (INT8/FP16), and energy efficiency, validating TwinLiteNet+ as a compelling solution for real-world autonomous driving systems. Code is available at https://github.com/chequanghuy/TwinLiteNetPlus.
@@ Abstract
Self-Driving Car , Lane segmentation , Drivable segmentation , BDD100K , Light-weight model , Embedded devices1 Introduction
The emergence of deep learning methodologies has driven significant growth in the field of autonomous vehicles, where visual perception plays a central role. Among various sensors, cameras remain the most cost-effective and offer high-resolution RGB data for scene understanding. Compared to LIDAR and Radar, which provide depth but are costly and color-insensitive, cameras offer a practical trade-off between cost computation and semantic detail. As a result, camera-based perception has become a dominant focus in the development of real-time driving systems, particularly those powered by convolutional neural networks.
In Advanced Driver Assistance Systems (ADAS) [2, 3, 4], camera-based panoptic driving perception serves as a critical component. By capturing comprehensive semantic information about the surrounding environment from images, vehicles can make optimal driving decisions. Prior studies [5, 6, 7, 8, 9, 10, 11, 12] have highlighted that segmentation plays a fundamental role in enabling ADAS to react promptly and accurately. Specifically, (1) Drivable area segmentation identifies regions where the vehicle can safely operate, thereby supporting precise control decisions. (2) Lane segmentation provides detailed insights into lane structure and positioning, assisting the vehicle in maintaining or changing lanes safely. While models are trained on large-scale datasets [13, 14, 15, 16, 17, 18] such as Cityscapes [19] offer detailed scene-level semantic labels, their granularity is often less relevant for immediate control decisions. In contrast, directly segmenting drivable regions and lanes provides more targeted and operationally helpful outputs for autonomous navigation.
Recent research has increasingly shifted from single-task models [5, 6, 7, 8, 9] to multi-task learning models [10, 11, 12], which aim to address multiple segmentation tasks concurrently. These models exploit shared encoder representations to reduce computational redundancy while maintaining task-specific decoder branches to preserve accuracy. However, achieving high performance often needs architectural complexity models, which may be impossible in real-world deployment. Despite being designed for real-time applications, many existing multi-task approaches [10, 20, 21, 22] are evaluated primarily on high-end GPUs, ignoring the strict constraints posed by embedded systems-such as limited processing power, energy efficiency, and memory capacity.
In this work, we introduce TwinLiteNet+, a model tailored for real-time operation with optimal power consumption and hardware resources, proficient in simultaneously segmenting lanes and drivable areas. This model demonstrates competitive precision compared to models with similar tasks. There are four principal contributions of our research:
1.
We introduce TwinLiteNet+, a lightweight and computation-efficient multi-task model designed for drivable area and lane segmentation. The encoder leverages a hybrid design of stride-based dilated convolutions and depthwise dilated separable convolutions to balance receptive field size and computational cost. In the decoder, we introduce Upper Convolution Block (UCB) and Upper Simple Block (USB) modules, which enhance upsampling efficiency without introducing heavy computation, making the architecture suitable for real-time inference.
2.
TwinLiteNet+ is implemented in four configurations-Nano, Small, Medium, and Large-ranging from just 0.03M parameters / 0.57 GFLOPs (Nano) to 1.94M parameters / 17.58 GFLOPs (Large). This range of configurations enables flexible deployment across devices with varying computational capabilities, from ultra-low-power embedded systems to higher-end edge devices.
3.
Experimental results on the BDD100K dataset show that TwinLiteNet achieves 92.9% mIoU for drivable area segmentation and 34.2% IoU for lane segmentation, outperforming state-of-the-art lightweight models while maintaining low model complexity.
4.
We deploy TwinLiteNet+ on real-world embedded devices such as NVIDIA Jetson TX2 and Jetson Xavier, demonstrating real-time inference performance with low latency and efficient energy consumption. Furthermore, we show that the model is highly robust under INT8 and FP16 quantization settings.
@@
The remainder of the paper is represented as follows: We evaluate relevant models in Section 2 to grasp the benefits and drawbacks in the tasks of Drivable Area Segmentation, and Lane Segmentation with Multi-task approaches. The proposed TwinLiteNet+ presents an architecture with methods to boost model performance in Section 3. In Section 4, we conduct experiments on the BDD100K dataset, and the results show that our TwinLiteNet+ models outperform on latency and power efficiency. In the final Section, we provide some conclusions and future development directions.
2 Related Work
2.1 Semantic Segmentation for autonomous driving
Semantic segmentation is a fundamental subfield in computer vision that aims to assign a class label to each pixel in an input image. Unlike tasks such as image classification or object detection, semantic segmentation provides fine-grained information by precisely identifying the shape and boundaries of each object within a scene. Research in this area has primarily focused on scene understanding [13, 14, 15, 16, 17, 18, 23], with models typically trained and evaluated on densely annotated datasets such as Cityscapes [19]. Modern segmentation models are often built upon deep learning architectures and further enhanced through the integration of attention modules [24, 25, 26, 27], which help the model focus on informative regions and improve spatial representation learning. Although these approaches have achieved high performance in modeling complex scene contexts, they are not well suited for real-time applications such as autonomous driving, where computational efficiency are critical. In such scenarios, a more practical approach is to focus on task-relevant functional regions-such as drivable areas and lane segmentation-to enable timely and reliable driving decisions.
2.1.1 Signle-task Approaches
Drivable area segmentation enables models to identify navigable regions of the road directly. Several existing approaches [8, 7] handle the drivable area as a single unified class, encompassing both direct and alternative drivable areas. In contrast, more advanced models [28, 29, 30, 31, 32, 33] have demonstrated the ability to distinguish between direct and alternative areas when trained on comprehensive datasets such as BDD100K, thereby offering a more understanding of the drivable space. However, segmenting the road into fine-grained components such as individual lanes-particularly differentiating between vertical lanes and parallel lanes (e.g., stop lines, crosswalks)-requires lane detection models. Parashar et al.[9] proposed Spatial CNN (SCNN) with a new neural network architecture by generalizing traditional deep layer-by-layer to slice-by-slice convolution in feature mapping, allowing information to propagate across pixels along rows and columns within a layer despite time-consuming. On the other hand, Enet-SAD [6] utilizes a self-attention-guided filter method to assist low-level feature maps in learning from high-level feature maps, or recently, road marking methods [34, 35, 36, 37] has also brought much attention from the community. A novel detector CLRerNet [38] along with a proposed LaneIoU metric are presented to improve the confidence score further.
Although previous research highlighted the initial success of models in the single-task domain, the advantage of leveraging a model to address multiple tasks has led to recent studies demonstrating that a multi-task strategy is more appropriate for real-world applications, particularly when deployed on devices with limited computing capabilities, such as embedded systems. This method allows models to strike a balance between efficiency and complexity while concurrently performing two tasks: segmentation of the driver’s area and lane segmentation. Motivated by these advantages, we propose a multi-task architecture with a shared encoder and dual decoders, targeting both drivable area and lane segmentation tasks simultaneously.
@@ 2.2 Multi-task Approaches
In recent years, multi-task learning has gained considerable attention in autonomous driving, particularly for jointly addressing segmentation-related tasks such as drivable area segmentation, lane segmentation, and scene parsing. By leveraging shared representations among these spatially correlated tasks, multi-task frameworks can improve model generalization while significantly reducing computational overhead-an essential consideration for real-time applications. The availability of large-scale datasets such as BDD100K [1] and NYUv2 [39] has further accelerated research in this direction, enabling the development and benchmarking of various segmentation-based multi-task models. Several studies extend to auxiliary segmentation tasks, combining semantic segmentation with scene classification [40, 31] or depth estimation [41, 42] to enhance environmental understanding. Others have adopted dual-task configurations, such as jointly performing drivable area and lane segmentation [12, 11, 10, 43, 16, 15, 44], which are critical for navigation and path planning.
Among prior multi-task segmentation approaches, some methods [15, 16] repurpose single-task segmentation backbones and adapt them to multi-task settings. At the same time, [43, 44] are initially designed for joint detection and segmentation, yet provide segmentation-only configurations. However, these models are not optimized explicitly for segmentation synergy, often resulting in limited performance for multi-task segmentation. In contrast, models like IALaneNet [10] explicitly focus on inter-task feature fusion using attention mechanisms at the task level. These methods demonstrate improved cooperation between segmentation branches but are limited by complex architectures and high computational costs, making them less suitable for real-time or embedded deployment. To address the limitations of practical deployment, lightweight models such as TwinLiteNet and BILane [12, 11] have been proposed. TwinLiteNet, with only 0.4M parameters, is optimized for real-time inference on embedded platforms, though its shallow decoder constrains performance in challenging lane scenarios. BILane [11] improves upon this with boundary-aware and interactive attention modules in a compact 1.4M-parameter design. However, its segmentation accuracy remains not as strong compared to TwinLiteNet despite using approximately 3.5 times more parameters and relying on higher input image resolution. These works clearly show a trade-off between segmentation accuracy and inference speed. This motivates the need for a multi-task segmentation model that can balance accuracy and efficiency.
Inspired by TwinLiteNet [12], the proposed TwinLiteNet+ introduces a more powerful yet efficient architecture by enhancing both the encoder and decoder. The decoder is improved with additional convolutional layers after transposed convolutions to reconstruct fine-grained features better, while the encoder is designed to optimize computational efficiency without compromising feature extraction capability. Moreover, TwinLiteNet+ comes in four scalable configurations, enabling flexible deployment across a wide range of hardware platforms-from high-performance GPUs to resource-constrained embedded systems.
@@ 2.3 Optimizing Deep Learning Deployment on Embedded Devices
In recent years, deploying deep learning models on embedded devices has attracted considerable attention within the artificial intelligence community. The computation on these devices is distinguished by its ability to execute processes directly on units that are both cost-effective and task-specific. Despite this advantage, such capabilities are constrained by significant limitations in terms of computational power and storage capacity, presenting substantial challenges for the deployment of deep learning models on these platforms. To address these limitations, several model compression techniques have been proposed, such as 8-bit quantization [45], pruning [46], and knowledge distillation [47]. Among these, quantization is widely favored for its simplicity and hardware-friendliness. It reduces model precision to lower bit widths (e.g., INT8 or FP16), significantly decreasing memory usage and inference latency. Post-training quantization (PTQ) allows models trained in 32-bit floating-point to be quantized without retraining, whereas quantization-aware training (QAT) simulates quantization effects during training and typically yields higher accuracy. In contrast, pruning requires careful exploration of pruning schedules and sensitivity analysis across layers, often requiring iterative retraining to restore accuracy. This process is time-consuming and highly dependent on manual tuning or costly search algorithms, and it may not be generalized well across architectures. Similarly, knowledge distillation introduces a complex teacher-student training pipeline where a compact model (student) is trained to mimic the output distribution of a larger pre-trained model (teacher). This multi-stage training not only adds overhead but also does not offer guarantees about interpretability or alignment with the original model’s decision boundaries.
Howard et al. introduce MobileNet [48], a class of efficient neural network architectures tailored for embedded and mobile vision applications, leveraging depthwise separable convolutions to reduce both parameter count and computational complexity significantly. Based upon this foundation, some studies have shown that depthwise dilated separable convolutions outperform standard dilated convolutions in classification tasks, achieving 67.9% accuracy with only 123M FLOPs-compared to 69.2% accuracy with 478M FLOPs using standard dilation-highlighting that the latter offers only marginal gains in accuracy at nearly 4× the computational cost. Beyond architectural improvements, numerous model compression and optimization strategies have been proposed for low-resource deployment [49, 50]. Frameworks such as TensorFlow Lite, TensorRT, NCNN, and MNN [51] have played a pivotal role in facilitating deployment by supporting quantization, operator fusion, and backend-specific accelerations.
Despite significant progress, achieving an optimal trade-off between accuracy, latency, and computational efficiency remains a persistent challenge-especially in real-time applications. While state-of-the-art models often attain impressive accuracy, their deployment on resource-constrained devices is limited by substantial latency and memory overhead. This trade-off is particularly critical in embedded and edge scenarios, where computational resources are inherently limited. In this study, we introduce a computation-efficient model specifically designed for deployment on such low-computation devices without resorting to architecture search techniques like pruning or additional training procedures such as knowledge distillation. Instead, the proposed model exhibits strong performance under low-precision inference settings, made possible through model compression via quantization (Sec. 4.5). Our architecture integrates a carefully optimized combination of dilated convolutions and depthwise separable convolutions within the encoder, balancing representational capacity and computational efficiency. This design not only preserves high segmentation accuracy but also achieves low inference latency, rendering it well-suited for real-time deployment in practical environments.